OpenStudy (anonymous):

The number of cartoons X, watched by Mrs. Kelly's first grade class on Saturday morning is shown below. X P(X) 0 0.15 1 0.20 2 0.30 3 0.10 4 0.20 5 0.05 Give the standard deviation for the probability distribution above.

OpenStudy (anonymous):

The standard deviation \[\sigma\] is the square root of the variance: \[\sigma = \sqrt{\mathbb{V}(X)} = \sqrt{\mathbb{E}(X^2) - (\mathbb{E}(X)^2})\]where \[\mathbb{E}(X)\]is the expected value of X. By definition, \[\mathbb{E}(X) = \sum_{x \in D} x p(x)\]where D is the values X may take (in this case \[D = \{0, 1, 2, 3, 4, 5\})\]and \[p(x) = \mathbb{P}(X = x)\]is the probability that X equals x. Then, \begin{eqnarray*}\mathbb{E}(X) &=& 0 \cdot p(0) + 1 \cdot p(1) + 2 \cdot p(2) + 3 \cdot p(3) + 4 \cdot p(4) + 5 \cdot p(5) \\ &=& 2.15. \end{eqnarray*}Similarly, \begin{eqnarray*}\mathbb{E}(X^2) &=& 0^2 \cdot p(0) + 1^2 \cdot p(1) + 2^2 \cdot p(2) + 3^2 \cdot p(3) + 4^2 \cdot p(4) + 5^2 \cdot p(5) \\ &=& 6.75, \end{eqnarray*}so \[\mathbb{V}(X) = 6.75 - 2.15 = 4.6\]and\[\sigma = \sqrt{4.6} \approx 2.144.\]